OpenStudy (anonymous):

what is the value of ∑_(k=1)^4▒〖(1/4)^k〗

OpenStudy (anonymous):

Missing No. ?!

OpenStudy (anonymous):

\[ \sum_{k=1}^4\left(\frac 14\right)^k = \left(\frac 14\right)^1+\left(\frac 14\right)^2+\left(\frac 14\right)^3+\left(\frac 14\right)^4 \]Is one way to do it.

OpenStudy (anonymous):

You could use \[ \sum_{k=0}^{n-1}r^k =\frac{1-r^n}{1-r} \]